OpenStudy (anonymous):

what is the point-slope equation of the line with the slope -10 that goes through the point (1,4)?

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given a point (x1, y1) so plug in m = slope = -10 and x1 = 1 , y1 = 4 into y - y1 = m(x - x1)